Evaluate the following integral.

∫(1 to 4)∫(0 to 4) (6x−3y)dxdy=
1
Expert's answer
2020-10-26T19:55:01-0400

The double integral is evaluated as,


"\int_{1}^{4}\int_{0}^{4}(6x-3y)dxdy" "=\int_{1}^{4}[6(\frac{x^2}{2})-3yx]_{0}^{4}dy"


"=\int_{1}^{4}[3(4^2-0^2)-3y(4-0)]dy"


"=\int_{1}^{4}(48-12y)dy"


"=[48y-12(\frac{y^2}{2})]_{1}^{4}"


"=[48(4-1)-6(4^2-1^2)]"


"=[48(3)-6(16-1)]"


"=144-6(15)"


"=144-90"


"=54"


Therefore, the double integral is "\int_{1}^{4}\int_{0}^{4}(6x-3y)dxdy=54" .